# Provably Secure Disambiguating Neural Linguistic Steganography

### v1

@@ IV-B3 Efficiency
Additionally, we conduct timing experiments on language models for the random sampling process, the ambiguity-unaware steganography generation process, and the disambiguating steganography generation process.
  • Ave Time: the average time consumed for embedding each bit of message, which is derived by dividing the total time of the entire process by the total length of the embedded message.
@@ IV-C Results and Analysis
IV-C1 Comparison
We compared the effectiveness of our method with that of token-removal-based methods in Discop for resolving ambiguity. The English and Chinese experimental results are presented in Table I and Table II, respectively. In these tables, “Discop” refers to the ambiguity-unaware provably secure steganographic algorithm that does not utilize any ambiguity resolution methods. “Discop +” refers to the performance after incorporating the disambiguation algorithm. “Basic” and “MWIS” correspond to the basic version and the version with MWIS of the token-removal-based disambiguating algorithm, respectively. “SyncPool” refers to the method that we proposed in this paper. Ave Time, Ave KLD, Max KLD, and Utilization are the key metrics for measuring the performance of steganography methods, and Ave PPL, Total Time, and Capacity, are also listed for reference. The experimental results are described and analyzed as follows.
Effectiveness
@@ IV-C1 Comparison
TABLE IV: The additional time consumption introduced by disambiguating algorithms
Model/Steganography k Ratio
+ Basic + MWIS + SyncPool
Llama2 Discop 16 1.00 1.02 1.01
32 1.01 1.04 1.02
64 1.02 1.09 1.02
128 1.05 2.04 1.01
256 1.19 7.72 1.01
Baichuan2 Discop 16 1.21 1.02 1.00
32 1.25 1.03 1.00
64 1.31 1.05 1.00
128 1.37 1.19 1.01
256 1.06 2.13 0.70
Time Efficiency
Compared to ambiguity-unaware embedding, our disambiguating provably secure steganography does not consume much additional time. Table IV shows the ratio of the time spent on provably secure steganography using different disambiguation methods to the time spent without incorporating them. During message embedding, SyncPool groups ambiguity pools, resulting in a distribution used by the steganographic embedding algorithm that has fewer elements compared to the original distribution, making embedding simpler and more efficient. At the same time, some additional time is consumed when running operations such as probability sorting, identifying ambiguous tokens, and synchronous sampling. The above operations incur much less additional time overhead compared to the process of finding MWISs in candidate pools.
@@ IV-D Ambiguity Frequency
TABLE V: The vocabulary sizes of different models, as well as the frequency of prefix relationships between tokens randomly sampled within different top-$k$ and other tokens.
@@
To further investigate the influence of segmentation ambiguity across various models, we conducted experiments on several common English and Chinese open-source models. Table V illustrates the vocabulary sizes of various models and the frequency of ambiguity under different top-$k$ truncation settings. Here, we randomly sample context $100$ times, with $100$ tokens sampled each time. The likelihood of segmentation ambiguity is represented by the proportion of tokens selected via random sampling in each step that share a prefix relationship with the other $k-1$ tokens. Overall, within the same language, as the number of tokens in the candidate pool increases, the frequency of ambiguity also tends to increase. Additionally, with smaller vocabulary sizes, the ambiguity frequency correspondingly increases. When comparing different languages, languages such as Chinese, which lack natural segmentation, have a greater probability of experiencing segmentation ambiguity than languages such as English.
@@ IV-E Examples of Stegotext
TABLE VI: Examples of generating stegotext in different languages using Discop [20] with and without disambiguation on Baichuan2. $k=256$.
@@
Table VI provides an example of employing Discop with and without our proposed disambiguating algorithms on Baichuan2 to continue writing 100 tokens under $k=256$ truncation. The semantics of stegotexts are all relevant to the given context.

### v2

@@ IV-B3 Efficiency
Additionally, we conduct timing experiments on language models for the ambiguity-unaware steganography generation process and the disambiguating steganography generation process. The time at which the model predicts the probability distribution is ruled out.
  • Ave Time: the average time consumed for embedding each bit of message, which is derived by dividing the total time of embedding by the total length of the embedded message.
@@ IV-C Results and Analysis
We compared the effectiveness of our method with those of token-removal-based methods for resolving ambiguity. The English and Chinese experimental results are presented in Table I and Table II, respectively. In these tables, “Discop” refers to the ambiguity-unaware provably secure steganographic algorithm that does not utilize any ambiguity resolution methods. “Basic” and “MWIS” correspond to the basic version and the version with MWIS of the token-removal-based disambiguating algorithm, respectively. “SyncPool” refers to the method that we proposed in this paper. The experimental results are described and analyzed as follows.
IV-C1 Effectiveness
@@ IV-C3 Utilization Efficiency
IV-C4 Time Efficiency
Fig. 4 shows the variation of embedding time (generation of $100$ tokens) with the value of $k$ for the steganographic embedding algorithm using different disambiguation methods. Compared to ambiguity-unaware embedding, our disambiguating provably secure steganography does not consume much additional time.
Figure 4: Time consumption introduced by disambiguating algorithms. (a) Llama2; (b) Baichuan2.
@@ IV-D Ambiguity Frequency
TABLE IV: The Vocabulary Sizes and the Frequency of Prefix Relationships of Different Models.
@@
To further investigate the influence of segmentation ambiguity across various models, we conducted experiments on several common English and Chinese open-source models, GPT-2444https://huggingface.co/openai-community/gpt2, Llama-2-7b, GPT-2-chinese [^5] https://huggingface.co/benjamin/gpt2-wechsel-chinese, and Baichuan2-7b. Table IV illustrates the vocabulary sizes of various models and the frequency of ambiguity under different top-$k$ truncation settings. Here, we randomly sample context $100$ times, with $100$ tokens sampled each time. The likelihood of segmentation ambiguity is represented by the proportion of tokens selected via random sampling in each step that share a prefix relationship with the other $k-1$ tokens. Overall, within the same language, as the number of tokens in the candidate pool increases, the frequency of ambiguity also tends to increase. Additionally, with smaller vocabulary sizes, the ambiguity frequency correspondingly increases. When comparing different languages, languages such as Chinese, which lack natural segmentation, have a greater probability of experiencing segmentation ambiguity than languages such as English.
@@ IV-E Examples of Stegotext
TABLE V: Examples of generating stegotext in different languages using Discop [20] with and without disambiguation on Baichuan2. $k=256$.
@@
Table V provides an example of employing Discop with and without our proposed disambiguating algorithms on Baichuan2 to continue writing 100 tokens under $k=256$ truncation. The semantics of stegotexts are all relevant to the given context.